Three point charges, q1=+5x10-6C, q2=-8x10-6C and q3=+2x10-6C, are placed on the corners of an equilateral triangle with sides of length 120 mm. Find the magnitude and direction of the net force over charge q1=+5x10-6C (use 2 decimal plates)
1
Expert's answer
2019-10-16T09:37:08-0400

"q_2" will attract "q_1" and "q_3" will repel "q_1"

Force on "q_1" by "q_3=F_1=\frac{K\times q_1\times q_3}{r^2}=6.250N"

Force on "q_1" by "q_2=F_2=\frac{K\times q_1\times q_2}{r^2}=25N"

Hence resultant force on "q_1=\sqrt{F_1^2+F_2^2+2F_1F_2cos\theta}=507.81N"